Net Present Value Method, Internal Rate of Return Method, and Analysis for a Service Company

The management of Advanced Alternative Power Inc. is considering two capital investment projects. The estimated net cash flows from each project are as follows:

Year Wind Turbines Biofuel Equipment
1 $340,000 $650,000
2 340,000 650,000
3 340,000 650,000
4 340,000 650,000

The wind turbines require an investment of $880,260, while the biofuel equipment requires an investment of $1,855,750. No residual value is expected from either project.

Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.353 2.991
6 4.917 4.355 4.111 3.785 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

Required:

1a. Compute the net present value for each project. Use a rate of 10% and the present value of an annuity of $1 in the table above. If required, use the minus sign to indicate a negative net present value. If required, round to the nearest whole dollar.

Wind Turbines Biofuel Equipment
Present value of annual net cash flows $ $
Less amount to be invested
Net present value $ $

1b. Compute a present value index for each project. If required, round your answers to two decimal places.

Present Value Index
Wind Turbines
Biofuel Equipment

2. Determine the internal rate of return for each project by (a) computing a present value factor for an annuity of $1 and (b) using the present value of an annuity of $1 in the table above. If required, round your present value factor answers to three decimal places and internal rate of return to the nearest whole percent.

Wind Turbines Biofuel Equipment
Present value factor for an annuity of $1
Internal rate of return % %
Answer #1
1a
Wind Turbines Biofuel Equipment
Present value of annual net cash flows 1077800 2060500
Less amount to be invested 880260 1855750
Net present value 197540 204750
1b
Present Value Index
Wind Turbines 1.22 =1077800/880260
Biofuel Equipment 1.11 =2060500/1855750
2
Wind Turbines Biofuel Equipment
Present value factor for an annuity of $1 2.589 2.855
Internal rate of return 20% 15%
Workings:
Wind Turbines Biofuel Equipment
Present value of annual net cash flows =340000*3.170 =650000*3.170
Wind Turbines Biofuel Equipment
Present value factor for an annuity of $1 =880260/340000 =1855750/650000